If A + B =1, B + C = 2, C + D = 3, ... X + Y = 25, Y + Z = 26, find A - Z
Answer by richard1234(7193)

Suppose we take the first equation, subtract the second equation, add the third, subtract the fourth, etc...in other words,

(A+B) - (B+C) + (C+D) - ... -(Y+Z).

This sum telescopes to A-Z, so we have A-Z = 1-2+3-4+...-26 = -13.